OpenStudy (anonymous):

Find the inverse of the function. f(x) = x3 - 3

OpenStudy (anonymous):

to find the inverse, let f(x) = y. y = x^3 - 3 now swap x and y. x = y^3 - 3 then put it back in the form y = f(x) y^3 = 3 + x \[y = \sqrt[3]{3 + x} = f^{-1}(x)\]

OpenStudy (anonymous):

thanks ya!